OpenStudy (vanz):

how do you solve this equation? x+3/2=8/x-3?

OpenStudy (saifoo.khan):

\[\frac{x+3}{2} = \frac{8}{x-3}\]Cross multiply, \[(x-3)(x+3)=8(2)\] Use foil to solve the Left hand side and so on. :)

hero (hero):

Did you find x yet vanz?

OpenStudy (vanz):

wait is the answer x=-5 , x=5 ???

OpenStudy (saifoo.khan):

right.

OpenStudy (vanz):

Thank you!! :D